OpenStudy (anonymous):

Simplify.square root 27/12 will give metal A.-3/2 B.2/3 C.3/2 D.3

pooja195 (pooja195):

Can that be simplified?

pooja195 (pooja195):

27/12

OpenStudy (anonymous):

yes

pooja195 (pooja195):

To what?

OpenStudy (astrophysics):

\[\sqrt{\frac{ 27 }{ 12 }} \implies \left( \frac{ 27 }{ 12 } \right)^{1/2} = \frac{ 27^{1/2} }{ 12^{1/2} }\]

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi misty

pooja195 (pooja195):

Wow that just looks like a mess now.

OpenStudy (astrophysics):

\[\implies \frac{ \sqrt{27} }{ \sqrt{12} } = \frac{ 3\sqrt{3} }{ 2\sqrt{3} }\]

OpenStudy (misty1212):

first step is to note that \[\frac{27}{12}=\frac{9}{4}\] then take teh square root of the top and bottom separately

OpenStudy (astrophysics):

Should get used to the exponent rules

OpenStudy (mertsj):

\[\sqrt{\frac{27}{12}}=\sqrt{\frac{9\times3}{4\times3}}=\sqrt{\frac{9}{4}}=\frac{\sqrt{9}}{\sqrt{4}}=\frac{3}{2}\]

OpenStudy (misty1212):

forget the fractional exponents, they do not help your cause at all reduce first, take the square root last

OpenStudy (astrophysics):

No, don't forget just know there's more then one way of doing it, and all you have to do is cancel out the sqrt(3)'s